# LUDVIG: Learning-free Uplifting of 2D Visual features to Gaussian Splatting scenes

### v4

@@ C.2 Visual comparisons of uplifted features
Figure 9: Comparison to GaussianEditors’s uplifting. Comparison of PCA visualization of uplifted features between LUDVIG’s and GaussianEditor’s aggregation (Chen et al., 2024).
C.4 Visualization of CLIP localization and segmentation results
@@
In this section, we present illustrations of the impact of the diffusion process (Figure 10), and comparative visualizations of localization heatmaps with LangSplat and LERF (Figure 11).
C.4.1 Impact of segmentation with SAM and DINOv2-guided graph diffusion
@@
Figure 10 shows 2D segmentation masks colored by CLIP relevancy scores, obtained with and without leveraging SAM and/or DINOv2-guided graph diffusion for refining 3D relevancy scores.
Direct segmentation from raw 3D relevancy scores.
@@
Isolating a specific object in the scene directly based on CLIP relevancy scores is challenging: the segmentation masks obtained by automatic thresholding include parts of other objects with similar features, like for the sheep (segmentation of the bear nose) and the spoon. The segmentation might also cover surroundings of the object of interest simply due to the low resolution of CLIP visual features, such as in the knife example.
2D segmentation with SAM.
@@
SAM delivers a precise 2D segmentation of the object covered by points with the highest relevancy scores. However, point prompts may not span the entire object, resulting in undersegmentation, like for the sheep. In some cases, point prompts with the highest relevancy may even be located on the wrong object, resulting in an entirely wrong segmentation (e.g., the bowl segmented instead of the spoon).
Relevancy score refinement with graph diffusion based on 3D DINOv2 features.
@@
The graph diffusion process starts with positive weights for Gaussians with the highest relevancy scores, and propagates their weight to neighbors with similar DINOv2 features. However in cases where the object of interest consists of multiple subparts (e.g. for the sheep), the final distribution of weights may be inhomogeneous and the automatic thresholding may select only one subpart. Also, if multiple close objects are to be segmented (e.g. with the knife), the final weights may cover surrounding Gaussians and the final thresholding might not clearly isolate the objects.
3D graph diffusion combined with 2D SAM segmentation.
@@
Combining 3D graph diffusion and 2D SAM segmentation helps solving the aforementioned problems observed when using either of the two approaches individually. The diffusion process allows selecting a large set of point prompts for SAM spanning the object of interest without covering other object with similar features, resulting in an accurate segmentation.
C.4.2 Qualitative comparison of open-vocabulary objet localization.
@@
Figure 11 illustrates open-vocabulary object localization with LERF (Kerr et al., 2023), LangSplat (Qin et al., 2024) and LUDVIG. Both LangSplat and LUDVIG correctly localize all four example objects. For queries such as the chopsticks, LangSplat’s localization is more precise, as the CLIP features are constructed by generating full image segmentation masks with SAM. This process is computationally expensive, as constructing a full segmentation mask requires querying SAM over a grid of points on the image and takes about 23s for a single image (on a GPU A6000 ADA), which amounts to an average of 80 minutes for a scene from the LERF dataset. However, it yields coherent instance-level CLIP representations, which is desirable for downstream segmentation tasks.
@@
Graph diffusionSAMGraph diffusionSAMRefer to captionRGB imageRefer to captionGround truthRefer to captionCLIP featuresRefer to captionDINOv2 featuresRefer to captionNo diffusion, no SAMRefer to captionDiffusion, no SAMRefer to captionNo diffusion, SAMRefer to captionDiffusion, SAM
Graph diffusionSAMGraph diffusionSAMGraph diffusionSAMRefer to captionRGB imageRefer to captionGround truthRefer to captionCLIP featuresRefer to captionDINOv2 featuresRefer to captionNo diffusion, no SAMRefer to captionDiffusion, no SAMRefer to captionNo diffusion, SAMRefer to captionDiffusion, SAM
Figure 10: Open-vocabulary object segmentation with and without using 3D graph diffusion (blue) and/or 2D SAM segmentation (green). Projections of 3D CLIP and DINOv2 features colored by three main PCA components and 2D segmentation masks colored by relevancy scores.
Figure 11: Qualitative comparisons of open-vocabulary 3D object localization on the LERF dataset. The red points are the model predictions and the black dashed bounding boxes denote the annotations. This figure is sourced and adapted from LangSplat’s website (https://langsplat.github.io/), licensed under a Creative Commons Attribution-ShareAlike 4.0 International License.

### v5

@@ D.4 Comparison to GaussianEditor’s uplifting
Figure G: Comparison to GaussianEditors’s uplifting. Comparison of PCA visualization of uplifted features between LUDVIG’s and GaussianEditor’s aggregation [6].
D.5 Visualization of CLIP segmentation results
@@
In this section, we present illustrations of the impact of the diffusion process (Fig. H), and comparative visualizations of localization heatmaps with LangSplat and LERF (Fig. I).
D.5.1 DINOv2-guided graph diffusion
@@
Fig. H shows 2D segmentation masks colored by CLIP relevancy scores, obtained with and without leveraging SAM and/or DINOv2-guided graph diffusion for refining 3D relevancy scores.
@@
Direct segmentation from raw 3D relevancy scores. Isolating a specific object in the scene directly based on CLIP relevancy scores is challenging: the segmentation masks obtained by automatic thresholding include parts of other objects with similar features, like for the sheep (segmentation of the bear nose) and the spoon. The segmentation might also cover surroundings of the object of interest simply due to the low resolution of CLIP visual features, such as in the knife example.
@@
2D segmentation with SAM. SAM delivers a precise 2D segmentation of the object covered by points with the highest relevancy scores. However, point prompts may not span the entire object, resulting in undersegmentation, like for the sheep. In some cases, point prompts with the highest relevancy may even be located on the wrong object, resulting in an entirely wrong segmentation (e.g., the bowl segmented instead of the spoon).
@@
Relevancy score refinement with graph diffusion based on 3D DINOv2 features. The graph diffusion process starts with positive weights for Gaussians with the highest relevancy scores, and propagates their weight to neighbors with similar DINOv2 features. However in cases where the object of interest consists of multiple subparts (e.g. for the sheep), the final distribution of weights may be inhomogeneous and the automatic thresholding may select only one subpart. Also, if multiple close objects are to be segmented (e.g. with the knife), the final weights may cover surrounding Gaussians and the final thresholding might not clearly isolate the objects.
@@
3D graph diffusion with 2D SAM segmentation. Combining 3D graph diffusion and 2D SAM segmentation helps solving the aforementioned problems observed when using either of the two approaches individually. The diffusion process allows selecting a large set of point prompts for SAM spanning the object of interest without covering other object with similar features, resulting in an accurate segmentation.
D.5.2 Comparisons for open-vocabulary localization
@@
Fig. I illustrates open-vocabulary object localization with LERF [19], LangSplat [35] and LUDVIG. Both LangSplat and LUDVIG correctly localize all four example objects. For queries such as the chopsticks, LangSplat’s localization is more precise, as the CLIP features are constructed by generating full image segmentation masks with SAM. This process is computationally expensive, as constructing a full segmentation mask requires querying SAM over a grid of points on the image and takes about 23s for a single image (on a GPU RTX 6000 ADA), which amounts to an average of 80 minutes for a scene from the LERF dataset. However, it yields coherent instance-level CLIP representations, which is desirable for downstream segmentation tasks.
@@
Figure H: Open-vocabulary object segmentation with and without using 3D graph diffusion (blue) and/or 2D SAM segmentation (green). Projections of 3D CLIP and DINOv2 features colored by three main PCA components and 2D segmentation masks colored by relevancy scores.
Figure I: Qualitative comparisons of open-vocabulary 3D object localization on the LERF dataset. The red points are the model predictions and the black dashed bounding boxes denote the annotations. This figure is taken and adapted from LangSplat’s website (https://langsplat.github.io/), licensed under a Creative Commons Attribution-ShareAlike 4.0 International License.